# Latent Equivariant Operators for Robust Object Recognition: Promise and Challenges

### v2

@@ Abstract
Despite the successes of deep learning in computer vision, difficulties persist in recognizing objects that have undergone group-symmetric transformations rarely seen during training—for example objects seen in unusual poses, scales, positions, or combinations thereof. Equivariant neural networks are a solution to the problem of generalizing across symmetric transformations, but require knowledge of transformations a priori. An alternative family of architectures proposes to learn equivariant operators in a latent space, from examples of symmetric transformations. Here, using simple datasets of rotated and translated noisy MNIST, we illustrate how such architectures can successfully be harnessed for out-of-distribution classification, thus overcoming the limitations of both traditional and equivariant networks. While conceptually enticing, we discuss challenges ahead on the path of scaling these architectures to more complex datasets.
@@ 1 Introduction
Several approaches have been proposed to increase robustness of deep networks to group transformations that are absent or only partially present in the training dataset. First, equivariant neural networks offer guarantees of robustness to target group transformations (Cohen et al., 2019; Bekkers, 2019). However, transformations need to be pre-specified mathematically in order to construct the relevant architecture for them. Second, data augmentations schemes, used in combination with either supervised or self-supervised learning objectives (e.g., Benton et al., 2020; Zbontar et al., 2021; Brehmer et al., 2024), can ensure some level of invariance to pre-specified transformations. However, for optimal results, transformations need to be sampled uniformly across the entire range of parameters seen at test time (Gerken and Kessel, 2024; Perin and Deny, 2025). This is not always possible, in particular when one is only given examples of transformations in a limited range. In a third line of work, methods have been proposed that learn group transformations from examples, hereafter referred to as latent equivariant operator methods (Culpepper and Olshausen, 2009; Sohl-Dickstein et al., 2017; Connor and Rozell, 2020; Dupont et al., 2020; Bouchacourt et al., 2021; Keller and Welling, 2021; LeCun, 2022; Connor et al., 2024). These methods learn an encoder jointly with a latent space operator, such that after training the latent space operator produces transformations that are approximately equivariant to the transformations present in the dataset. These methods offer a promising alternative for out-of-distribution object recognition. [^1] Latent equivariant operator methods should be distinguished from disentanglement methods; the latter can be seen as special cases of the former, where the latent operator is confined to a subspace (Higgins et al., 2018)—but see topological defects induced by subspace disentanglement in Bouchacourt et al. (2021).
@@
Here, we illustrate how equivariant latent operator methods can successfully be applied to out-of-distribution classification problems. For simplicity, we strip down the complexity of the datasets and models to their minimal components necessary to establish the superiority of these methods over others. On rotated and translated noisy MNIST, we train a neural encoder jointly with a latent operator to learn equivariance over a limited range of transformations. At the same time, we train a classifier, taking as input the latent space, on the same range of transformations. At inference time, we test our model over a range of transformations that were not seen during training, assessing its extrapolation and composition abilities. We use a K-nearest-neighbour strategy to select the operator action most likely to revert the object back to its canonical pose. We establish the superiority of our method over comparable architectures on classifying inputs outside the training range. We conclude with a discussion on the challenges that are expected when scaling these methods to more complex datasets.
@@ 2 Methods
We use a simple feed-forward architecture. The encoder consists of a single linear layer that maps the flattened input into a latent representation to be transformed with the shift operator. When there are a combination of transformations, we use stacked encoders and operators for each transformation. We set the latent dimensionality to 70 to accommodate the orders of the transformation groups considered. The pre-defined latent operator follows the discrete construction of Bouchacourt et al. (2021): each block has a size equal to the order of the corresponding transformation group and is repeated along the diagonal to match the latent dimensionality. To evaluate whether equivariance can be learned from scratch, we also consider a learnable operator variant initialized as the Q factor of a QR decomposition (to ensure orthogonality) of a randomly sampled matrix. A classifier, stacked after the encoder, is a two-layer MLP with a sigmoid nonlinearity in the hidden layer, producing class logits from the transformed latent features. Further details are provided in Appendix A.
@@ 4 Discussion
Here we demonstrated, in a minimal setup, how latent equivariant operators can be leveraged to classify out-of-distribution samples—by means of extrapolation and combination of symmetric transformations. Unlike in equivariant neural networks, the symmetries did not need to be mathematically pre-specified. Unlike in data augmentation schemes, the method did not require coverage of the full range of transformation parameters during training. Intuitively, this is because the recursive application of operators in latent space allows the extrapolation of transformations beyond the training range, and in unseen combinations. Latent equivariant operator methods thus offer a promising avenue for robust, human-like [^2] Note that latent operators can be conceived as performing internal changes of perspective akin to mental simulation in humans (e.g., Shepard and Metzler, 1971; Khazoum et al., 2025). object recognition, and warrant further investigation.
To date, latent equivariant operator methods have not been shown to scale to large and complex datasets in the out-of-distribution domain. We identify a number of theoretical and implementation challenges ahead. First, we do not know theoretically the certainty with which we can expect operators to remain equivariant beyond the training range of transformation parameters. Empirically, we find classification performance to degrade somewhat outside the training range. Moreover, we do not know at what layer of an architecture such operators should be situated. For the affine transformations explored in this minimal study, a single linear layer suffices to recast the representation in a way suitable to the operator (whether learned or fixed), as described in previous theory (Bouchacourt et al., 2021). For more complex transformations, for example transformations happening in a latent space not directly accessible from pixel space (e.g., in-depth 3D rotation), it is unclear how many layers would be required. Empirically, latent equivariant operators have been shown to successfully emulate in-depth 3D rotations (Dupont et al., 2020). Finally, the choice of the functional form of latent operators is another open question, involving fundamental notions in representation theory and topology. The resolution of these theoretical questions should provide guidance for the successful implementation of latent equivariant operator approaches at scale.
@@ Acknowledgments
We thank members of the BRAIN lab at Aalto University, Ivan Vujaklija, Mansour Taleshi, and T. Andy Keller for useful discussions. This work was supported in part by the Academy of Finland under Grant 3357590 for S.D, and by the Helsinki Institute for Information Technology (HIIT) under Contract No. 9125064HIIT for M.D.T.
@@ A.3 Learned Operator
The hard-coded constructions presented in Section A.1 demonstrate the existence of a valid operator that satisfies the desired transformation behavior for this task. While these fixed operators provide a constructive proof of feasibility, they are not necessarily optimal when integrated into a larger learning pipeline. To allow the operator to adapt to the data and interact with other learned components, we parameterize the operator and optimize it jointly with the rest of the model.

### v4

# Latent Equivariant Operators for Robust Object Recognition: Promises and Challenges

@@ Abstract
Despite the successes of deep learning in computer vision, difficulties persist in recognizing objects that have undergone group-symmetric transformations rarely seen during training—for example objects seen in unusual poses, scales, positions, or combinations thereof. Equivariant neural networks are a solution to the problem of generalizing across symmetric transformations, but require knowledge of transformations a priori. An alternative family of architectures proposes to learn equivariant operators in a latent space, from examples of symmetric transformations. Here, using simple datasets of rotated and translated noisy MNIST, we illustrate how such architectures can successfully be harnessed for out-of-distribution classification, thus overcoming the limitations of both traditional and equivariant networks. While conceptually enticing, we discuss challenges ahead on the path of scaling these architectures to more complex datasets. Our code is available at https://github.com/BRAIN-Aalto/equivariant_operator.
@@ 1 Introduction
Several approaches have been proposed to increase robustness of deep networks to group transformations that are absent or only partially present in the training dataset. First, equivariant neural networks offer guarantees of robustness to target group transformations (Cohen et al., 2019; Bekkers, 2019). However, such approaches require complete a priori knowledge of the transformation. In particular, the transformation group structure (e.g., cyclic group of a certain order) and its specific representation (e.g., rotations or translations) must be specified mathematically to construct the relevant architecture. Second, data augmentations schemes, used in combination with either supervised or self-supervised learning objectives (e.g., Benton et al., 2020; Zbontar et al., 2021; Brehmer et al., 2024), can ensure some level of invariance to pre-specified transformations. However, for optimal results, transformations need to be sampled uniformly across the entire range of parameters seen at test time (Gerken and Kessel, 2024; Perin and Deny, 2025). This is not always possible, in particular when one is only given examples of transformations in a limited range. In a third line of work, methods have been proposed that learn group transformations from examples, hereafter referred to as latent equivariant operator methods (Culpepper and Olshausen, 2009; Sohl-Dickstein et al., 2017; Connor and Rozell, 2020; Dupont et al., 2020; Bouchacourt et al., 2021; Keller and Welling, 2021; LeCun, 2022; Connor et al., 2024). These methods learn an encoder jointly with a latent space operator, such that after training the latent space operator produces transformations that are approximately equivariant to the transformations present in the dataset. These methods offer a promising alternative for out-of-distribution object recognition. [^1] Latent equivariant operator methods should be distinguished from disentanglement methods; the latter can be seen as special cases of the former, where the latent operator is confined to a subspace (Higgins et al., 2018)—but see topological defects induced by subspace disentanglement in Bouchacourt et al. (2021).
@@
Here, we illustrate how equivariant latent operator methods can successfully be applied to out-of-distribution classification problems. For simplicity, we strip down the complexity of the datasets and models to their minimal components necessary to establish the superiority of these methods over others. On rotated and translated noisy MNIST, we train a neural encoder jointly with a latent operator to learn equivariance over a limited range of transformations. At the same time, we train a classifier taking as input the latent space, on the same range of transformations. At inference time, we test our model over a range of transformations that were not seen during training, assessing its extrapolation and composition abilities. We use a K-nearest-neighbour strategy to select the operator action most likely to revert the object back to its canonical pose. We establish the superiority of our method over comparable architectures on classifying input samples outside the training range. Although our work builds on previous work—e.g., Bouchacourt et al. (2021); Connor and Rozell (2020)—we extend this prior work in several ways: (1) we demonstrate that latent equivariant operator methods can be used for classification outside the range of transformations seen during training, (2) while not specifying transformation parameters at test time, (3) and with a learnable operator necessitating only the specification of a weak periodicity prior. We conclude with a discussion on the challenges ahead on the path of scaling these methods to real world images and more complex datasets.
@@ 2 Methods
We use a simple feed-forward architecture. The encoder consists of a single linear layer that maps the flattened input into a latent representation to be transformed with the shift operator. When there are a combination of transformations, we use stacked encoders and operators for each transformation. We set the latent dimensionality to 70 to accommodate the orders of the transformation groups considered. The pre-defined latent operator follows the discrete construction of Bouchacourt et al. (2021): each block is a shift matrix with a size equal to the order of the corresponding transformation group and is repeated along the diagonal to match the latent dimensionality. To evaluate whether equivariance can be learned from scratch, we also consider a learnable operator variant initialized as the orthogonal factor $Q$ of a QR decomposition [^2] QR initialization offers a stable orthogonal starting point for optimization; simpler initializations such as zero or unconstrained random matrices led to less reliable convergence in our experiments. of a randomly sampled matrix. A classifier, stacked after the encoder, is a two-layer MLP with a sigmoid nonlinearity in the hidden layer, producing class logits from the transformed latent features. Further details are provided in Appendix A.
@@ 4 Discussion
Here we demonstrated, in a minimal setup, that latent equivariant operators can be leveraged to classify out-of-distribution samples—by means of extrapolation and combination of symmetric transformations. Unlike in equivariant neural networks, the symmetries did not need to be mathematically pre-specified. Unlike in data augmentation schemes, the method did not require coverage of the full range of transformation parameters during training. Intuitively, this is because the recursive application of operators in latent space allows the extrapolation of transformations beyond the training range, and in unseen combinations. Furthermore, the method does not require knowledge of the transformation parameter at inference time, as the latent transformation can be estimated through canonicalization in representation space. Latent equivariant operator methods thus offer a promising avenue for robust, human-like [^3] Note that latent operators can be conceived as performing internal changes of perspective akin to mental simulation in humans (e.g., Shepard and Metzler, 1971; Khazoum et al., 2025). object recognition, and warrant further investigation.
To date, latent equivariant operator methods have not been shown to scale to large and complex datasets in the out-of-distribution domain. In this work, our experiments use a minimal controlled setup (MNIST with synthetic noisy backgrounds) and focus on invariant classification. Evaluating these methods on more realistic noise sources, broader transformation families, and more complex datasets, therefore, remains an important direction for future work. At the same time, several theoretical and implementation challenges remain. First, we do not know theoretically the certainty with which we can expect operators to remain equivariant beyond the training range of transformation parameters. Empirically, we find classification performance to degrade somewhat outside the training range. Second, we do not know at what layer of an architecture such operators should be situated. For the affine transformations explored in this minimal study, a single linear layer suffices to recast the representation in a way suitable to the operator (whether learned or fixed), as described in previous theory (Bouchacourt et al., 2021). For more complex transformations, for example transformations happening in a latent space not directly accessible from pixel space (e.g., in-depth 3D rotation), it is unclear how many layers would be required. Empirically, latent equivariant operators have been shown to successfully emulate in-depth 3D rotations (Dupont et al., 2020). Finally, the choice of the functional form of latent operators is another open question, involving fundamental notions in representation theory and topology. The resolution of these theoretical questions should provide guidance for the successful implementation of latent equivariant operator approaches at scale.
@@ Acknowledgments
We thank members of the BRAIN lab at Aalto University, Ivan Vujaklija, Mansour Taleshi for useful discussions, and Fabio Anselmi and T. Andy Keller for helpful comments on the manuscript. This work was supported in part by the Academy of Finland under Grant 3357590 for S.D, and by the Helsinki Institute for Information Technology (HIIT) under Contract No. 9125064HIIT for M.D.T.
@@ A.3 Learned Operator
The hard-coded constructions presented in Section A.1 demonstrate the existence of a valid operator that satisfies the desired transformation behavior for this task. While these fixed operators provide a constructive proof of feasibility, they are not necessarily optimal when integrated into a larger learning pipeline. To allow the operator to adapt to the data and interact with other learned components, we parameterize the operator and optimize it jointly with the rest of the model. In practice, the operator is initialized using the orthogonal factor of a QR decomposition of a random matrix to provide a stable starting point for optimization.